I am given the equation
y=-2x
The instructions I am given are as stated, Use a table of values and graph six sets of ordered pairs for the equation. Describe the graph.

Answer by Fombitz(32388)

You get to choose the ordered pairs.
Pick 6 x values.
Generate 6 y values.
Use those to plot the graph of the equations.
Let's choose x=-3,-2,-1,0,1,2,3 just to make it equal on both sides.
f(-3)=-2(-3)=6, (-3,6)
f(-2)=-2(-2)=4, (-2,4)
f(-1)=-2(-1)=2, (-1,2)
f(0)=-2(0)=0, (0,0)
f(1)=-2(1)=-2, (1,-2)
f(2)=-2(2)=-4, (2,-4)
f(3)=-2(3)=-6, (3,-6)
.
.
.
Now plot those ordered pairs.

Connect the points.
.
.
.

Looks like a straight line through the origin sloping downward.